OpenStudy (anonymous):

Use Part 1 of the Fundamental Theorem of Calculus to find the derivative of the function. g(s) = integral of s to 2 of (t − t^9)^3dt

OpenStudy (anonymous):

\[\int_s^2(t-t^9)^3dt=-\int_2^s(t-t^9)^3dt\] is a start then since the derivative of the integral is the integrand, replace all \(t\) by \(s\) and you are dond

OpenStudy (anonymous):

*done

OpenStudy (anonymous):

\[g(s)=\int\limits_{2}^{s}(t-t^9)^3dt\]

OpenStudy (anonymous):

that was the starting equation

OpenStudy (anonymous):

so you're saying insert S everywhere t is? then what?

OpenStudy (anonymous):

then you are done

OpenStudy (anonymous):

what about the 2 in the integral?

OpenStudy (anonymous):

the derivative of the integral is the integrand, that is all the 2 determines the constant, but the derivative of a constant is zero, so it doesn't change anything

OpenStudy (anonymous):

ohhhh you're right! i didn't catch that part! thanks alot satellite!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if you don't mind... can you explain to me the difference between integrand and integral? are they essentially the same thing?

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

\[\huge \int\limits_{s}^{2}\color{red}{(t-t^9)^3}dx\] The whole expression is an integral. Red part is the integrand.

OpenStudy (anonymous):

oh word! that makes sense... so the derivative of an integral is just the integrand...

terenzreignz (terenzreignz):

Much like how... \[\huge \sqrt{\color{red}{x^2+6x+9}}\] the whole expression is a radical, but the red part is called the radicand.

OpenStudy (anonymous):

wow that makes a lot of sense actually

terenzreignz (terenzreignz):

Derivative of an integral is the integrand? That makes sense if it were an indefinite integral, or a definite integral of the form... \[\huge \int\limits_{a}^{x}f(t) dt\] Then certainly, \[\huge \frac{d}{dx} \int\limits_{a}^{x}f(t) dt = f(x)\]

OpenStudy (anonymous):

so that's what happened with my problem right... change it to s and the derivative of the integral was the integrand of f(s)

terenzreignz (terenzreignz):

Yes. Exactly :)

OpenStudy (anonymous):

omg i feel so much smarter!

terenzreignz (terenzreignz):

Except, in your case, it s was the lower limit instead of the upper limit... in that case, since... \[\huge \int\limits_{s}^{2}f(t) dt=-\int\limits_{2}^{s}f(t) dt\] then... \[\huge \frac{d}{ds}\int\limits_{s}^{2}f(t) dt=-f(s)\]

OpenStudy (anonymous):

oh okay! makes sense. thanks terenzreignz